olivia7026:

Suppose that the number of cars manufactured at an automobile plant varies jointly as the number of workers and the time they work. If 85 workers can produce 136 cars in 8 hours, find the number of workers that can produce 175 cars in 7 hours.

Tutorial:

the appropriate response is 66 vehicles. 160 specialists - 32 vehicles - >2 hours 220 specialists - > x vehicles - > 3 hours Assuming the quantity of laborers rise, we anticipate that the number of cars should rise aswel, so they're straightforwardly corresponding. In the event that the quantity of hours worked increment we can anticipate that the number of cars should increment aswell, so they're straightforwardly relative. We would now be able to set the parts: (160*2)/(220*3) = 32/x 320/660 = 32/x x = (32*660/320) = 66 vehicles They'll fabricate 66 vehicles in those conditions

Tutorial:

My source is Brainly

Tutorial:

please close your post when you are done with it.

Tutorial:

@dude requesting for this post to be closed.